Two planes left St. Louis for Los Angeles at the same time. After 4 hours they were 700 miles apart. The slower plane traveled at 350 miles/hour. What was the speed of the faster plane?
At first, I used the formula:
350(4)+4x=700
I kept getting a negative answer and the answer is supposedly 525 miles/hour. Any help would be appreciated! I've been working on this problem for a good 45 minutes already so please help!

Answer by askmemath(368)

The slower plane travelled at 350mph for 4 hours.Which means that in 4 hours it covered 350x4 = 1400m
But the distance between the faster and slower plane is 700 miles.Which means the faster plane travelled a distance of 1400+700 = 2100miles
Total time the faster plane took to travel this distance is also 4 hours
So Speed = 2100/4 = 525miles/hour.